Calculate the % of radiant energy converted into kinetic energy of x – atoms. x – atom forms after dissociation of X2 molecules by the absorption of 5 quantum of radiation having wave length 5000 Å. If bond energy of X2 molecules is 250 Kj mol-1.
 
The answer is 79.10%. I want to know the process of solution.

Grade:11

2 Answers

Vikas TU
7 years ago
Calculate first Radiant energy from hc/k for k = 5000 A.
miltoply it with Na for energy per mole.
K.E = radiant energy – diss. energy
% = K.E./radint energy * 100
Sam Dwayne Johnson
7 years ago
Radiant energy= hc/L
=6.626x10-34 x 3x10^8/ 5000 A
=3.97x10-19 J
Joule/mol=3.97x10-19x6.022x10^23
=23.9x104 J/mol
Energy of 5 quanta=23.9x104 J/molx5
=1195KJ/mol
 
K.E.=ER-ED
=1195-250
=945 Kj/mol
 
Percentage=KE/ER X 100
=945/1195 X 100
=79.07%
or 79.1%
Hope this helped u
ATB